OpenStudy (anonymous):

A fence 6 feet tall runs parallel to a tall building at a distance of 7 feet from the building. What is the length of the shortest ladder that will reach from the ground over the fence to the wall of the building?

OpenStudy (mathmath333):

this has long derivation answer but has very nice formula to it u can find the derivation here http://archives.math.utk.edu/visual.calculus/3/applications.4/ here \(\Large q= height~~of~~fence\\ \Large p=distance~~between~~fence~~and~~building\) \(\Large I(a)=\sqrt{(q+p^{\frac{2}{3}}q^{\frac{1}{3}})^2+(p+p^{\frac{1}{3}}q^{\frac{2}{3}})^2)}\\ \Large =\sqrt{(6+7^{\frac{2}{3}}6^{\frac{1}{3}})^2+(7+7^{\frac{1}{3}}6^{\frac{2}{3}})^2)}\\ \Large =18.36\)

OpenStudy (anonymous):

Thanks so much!

OpenStudy (mathmath333):

yw